### v1

# Kernel Debiased Plug-in Estimation

Abstract
We consider the problem of estimating a scalar target parameter in the presence of nuisance parameters. Replacing the unknown nuisance parameter with a nonparametric estimator, e.g., a machine learning (ML) model, is convenient but has shown to be inefficient due to large biases. Modern methods, such as the targeted minimum loss-based estimation (TMLE) and double machine learning (DML), achieve optimal performance under flexible assumptions by harnessing ML estimates while mitigating the plug-in bias. To avoid a sub-optimal bias-variance trade-off, these methods perform a debiasing step of the plug-in pre-estimate. Existing debiasing methods require the influence function (IF) of the target parameter as input. However, deriving the IF requires specialized expertise and thus obstructs the adaptation of these methods by practitioners. We propose a novel way to debias plug-in estimators which (i) is efficient, (ii) does not require the IF to be implemented, (iii) is computationally tractable, and therefore can be readily adapted to new estimation problems and automated without analytic derivations by the user. We build on the TMLE framework and update a plug-in estimate with a regularized likelihood maximization step over a nonparametric model constructed with a reproducing kernel Hilbert space (RKHS), producing an efficient plug-in estimate for any regular target parameter. Our method, thus, offers the efficiency of competing debiasing techniques without sacrificing the utility of the plug-in approach.
1 Introduction
In recent decades, the field of statistical learning has witnessed a wave of monumental advancement in causal inference methodology. Spearheaded by the development of data-adaptive machine learning methods and efficiency theory, modern estimation methods achieve optimal performance under no, or minimal, assumptions on the true data-generating process (DGP). The statistical learning advancements grounded in non/semi-parametrics complement the available knowledge regarding the DGP under study and the corresponding scientific question. In particular, they provide formal statistical language which reflects uncertainty in both the background knowledge and the available data in order to draw reliable conclusions from statistical analyses. The development and usage of modern causal inference methods have great potential to advance scientific discovery and enhance reproducibility.
We make an important distinction between causal and statistical steps, which combined form a backbone of modern causal inference [26, 27, 9]. In the first step, one postulates a causal model, which leads to numerous strategies for the identification of a target parameter (estimand) [44, 39, 42]. In the second step, which is the focus of our work, one aims to estimate the target parameter without imposing any assumptions on the DGP outside the available knowledge about the process (e.g., known randomization probabilities). As part of the second step, most modern causal inference methods use techniques that rely on asymptotic analyses and semiparametric efficiency theory of estimation [4, 25, 50]. The considered target parameters are primarily pathwise-differentiable, making them potentially estimable at $\sqrt{n}$-rate, despite their dependence on nonparametric nuisance estimands. For pathwise-differentiable target parameters, the first-order bias can conveniently be described as the expectation of a function of the observed data and the nuisance parameters [19]. The resultant estimators based on modern causal inference methods are efficient (i.e., lowest possible asymptotic variance) within the class of regular and asymptotically linear estimators which allow for the construction of asymptotically valid confidence intervals and hypothesis tests [52, 54]. By contrast, a naive estimator based on substituting any machine learning estimates for nuisance parameters would yield a non-negligible first-order bias and a slower than $\sqrt{n}$-rate of convergence, leading to invalid asymptotic approximation and inferences.
The statistical concept of asymptotic linearity roughly corresponds to the ability to approximate the difference between an estimate and the true value of the target parameter as an average of independent and identically distributed random variables. If an estimator admits to such an asymptotically linear form, then obtaining the sampling distribution of the estimator is straightforward. More technically, we write the first order expansion of an asymptotically linear estimator as $\hat{\psi}-\psi\approx\sum_{i=1}^{n}\tilde{\psi}(O_{i})/n,$ where $\hat{\psi}$ and $\psi$ are the estimate and the truth respectively, while $\tilde{\psi}(O_{i})$ is the the corresponding influence function (IF) evaluated over $n$ samples of the random variable $O$ [4, 40, 25]. For a given pathwise-differentiable target parameter, there exists a set of possible estimators, each with a corresponding IF. The efficient influence function (EIF) is $\tilde{\psi}$ with the lowest sampling variance. Any estimator whose IF corresponds to the EIF is considered efficient, as it attains the Cramér–Rao lower bound.
While asymptotic linearity ensures an estimator has a tractable asymptotic distribution, it does not guarantee that it behaves nicely across a range of possible data-generating processes. A regular estimator attains the same limiting distribution even under small (of the order $1/\sqrt{n}$) perturbations of the true data-generating distribution, thereby enforcing robustness to distributional shifts. We restrict our attention to the class of RAL estimators in this work, owing to their numerous statistical advantages: robustness to distributional shifts and easy characterization of the sampling distribution. We emphasize that the most efficient regular estimator is theoretically guaranteed to be asymptotically linear [53]. Non-regular super-efficient estimators are possible, but they are more efficient on an infinitesimally small part of the model space and lack a stable asymptotic sampling distribution [1].
With the IF in hand, there are a few general techniques for constructing efficient asymptotically linear estimators of the target parameter: estimating equation (and double machine learning), bias correction (known as the one-step), and the targeted minimum loss-based estimator (TMLE). Building on the advancements in semiparametric statistics, Robins et al. [43] propose a general estimating equation approach that solves for the target parameter by setting the score equations to zero [38, 4]. Chernozhukov et al. [8] introduce double machine learning (DML), which characterizes the estimand as the solution to a population score equation. In DML, first-order bias is defined with respect to a property called Neyman orthogonality, which ensures that the score equations that define the estimand are first-order invariant to small changes in the nuisance parameters. Bickel et al. [4] propose a bias correction method, which adds the empirical average of the IF to an initial estimator. The most modern approach, TMLE, constructs a plug-in estimator of the target parameter by fluctuating the initial estimate in the direction of the IF; by doing so, it forces the nuisance estimates to solve a set of user-defined estimating equations [29, 26, 27]. While all approaches have the same asymptotic properties, they differ in finite samples. For example, by the shared quality of being a plug-in estimator, TMLE naturally respects the bounds imposed by the statistical model.
We emphasize that all commonly used techniques for building an efficient asymptotically linear estimator depend on the knowledge of the IF. These functions result from the mathematical study of the estimand and are often difficult to analytically derive and/or compute [16, 19]. In this paper, we propose a novel approach for constructing efficient estimators that builds upon the TMLE framework and a novel application of reproducing kernel Hilbert spaces (RKHS), which does not rely on the knowledge of the IF. More specifically, the proposed method eliminates the need for calculating the IF of the target parameter, thereby simplifying and allowing automation of the estimation process. In addition, it results in a single estimate of the DGP which is efficient for many target parameters, instead of having a different estimate for each new estimand. Our method combines the utility of plug-in estimators and all the favorable theoretical properties of the TMLE, resulting in asymptotically normal and efficient estimates of all possible pathwise-differentiable target parameters, given mild regularity conditions. We support these findings through numerical validations on various target parameters, including the average treatment effect, relative risk, and odds ratio. Our results indicate that the proposed estimator performs comparably to those described by existing works, highlighting its effectiveness in practical applications. In summary, our key contributions are as follows:
  • We construct a debiased distribution estimator $P^{\infty}_{n}$ that works as an efficient plug-in $\psi(P^{\infty}_{n})$ for any pathwise differentiable target parameter $\psi$.
  • Our estimator $P^{\infty}_{n}$ does not require the IF $\tilde{\psi}_{P}$ to be computed, and, unlike other popular approaches in the literature, the same $P^{\infty}_{n}$ can be used for any target parameter $\psi(P)$.
  • Combined with a bootstrap scheme, our estimation procedure offers a fully automated framework for estimation and inference of any pathwise differentiable parameter without requiring analytic derivations of the influence function, debiased/orthogonalized estimating equations, or efficiency bounds from the user.
  • Leveraging the universal approximation property of reproducing kernel Hilbert spaces, we show that under regularity conditions our proposed algorithm is guaranteed to perform as well as more targeted methods and validate these claims with empirical simulations.
1.1 Other Related Works
Computerized approaches to debiasing.
Several papers have considered the problem of numerically approximating the values of the IF via finite-differencing [17, 6]. Jordan et al [18] acknowledge that adaptation of existing debiasing approaches is hindered by the necessity of deriving the IF. To address this issue, they consider combining a numerical scheme for evaluating the IF with existing debiasing methods, aiming to produce an automated framework for constructing a debiased estimator of target parameters. By contrast, our proposed debiasing scheme, by construction, does not require the IF and can be applied with any pathwise differentiable parameter.
Undersmoothed Highly Adaptive Lasso.
This sieve-based maximum likelihood estimator (MLE) results in an asymptotically efficient substitution (i.e., plug-in) estimator even for large models [47]. This approach requires a special type of sieve that undersmooths the nuisance parameter by effectively over-fitting the data. A modern approach, the undersmoothed Highly Adaptive Lasso- (HAL-) MLE, is defined as the loss-based minimizer over a special type of functions that can be well approximated by a linear combination of tensor products of univariate spline-basis functions with a bounded L1 norm [3, 23, 22]. This particular sieve-based MLE, therefore, relies on the assumption that the IF is a cadlag function with a finite sectional variation norm, and the sieves are indexed based on a bound on the L1 norm. Practically, HAL constructs basis functions that grow exponentially with the number of covariates $d$ and polynomially with the number of samples $n$, and puts a bound on the L1 norm of the coefficients: this makes it computationally extremely inefficient in large models. Our approach, while having a similar goal, uses only $n$ basis functions for our procedure and does not depend on the covariate dimension, $d$. We emphasize that our proposed method also puts no explicit restrictions on the regression surface of the models that we consider.
Reproducing kernel Hilbert space methods.
RKHS methods [55, ch 12] have emerged as a powerful tool for analysis, modeling and computation in various fields, including machine learning, statistics, as well as applied and pure mathematics. To every RKHS there is associated a (reproducing) kernel, a positive-definite function that encodes the inner product and all functions within the function space. In machine learning, RKHS methods have gained popularity due to their ability to handle high-dimensional data and nonlinear relationships while allowing for efficient computation and tractable theory. Machine learning algorithms that leverage computational properties of RKHS that are also central to our work include kernel ridge regression [46] and other regularized estimators. Recently, kernel-based debiasing techniques [15] and nonparametric modeling of probability distributions [13, 45, 31, 35] have received significant attention. There are also direct debiased kernel methods [48, 10] and kernel methods for efficient estimation of target parameters that obtain stronger than semiparametric efficiency results under stronger assumptions [34].
2 Problem Setup and Preliminaries
Set up.
Let $O_{1},\ldots,O_{n}$ denote i.i.d. observations drawn from a probability distribution $P^{*}$ on the sample space $\mathcal{O}$. The distribution $P^{*}$ is unknown but is assumed to belong to a nonparametric collection $\mathcal{M}$, which consists of distributions on $\mathcal{O}$. Let $\psi:\mathcal{M}\rightarrow\mathbb{R}^{m}$ be a functional of the model $\mathcal{M}$, also referred to as the target parameter. We consider problems where the target parameter is jointly defined together with nuisance parameters, and will take as the nuisance the entire distribution $P^{*}$ following [29]. Our goal is to estimate $\psi(P^{*})$ on the basis of the $n$ observations. This will be accomplished with a plug-in estimator, i.e., by estimating $P^{*}$ and evaluating $\psi$ on this estimate. Since our method produces a single debiased estimate $\widehat{P}$ that does not depend on $\psi$, we will assume that $m=1$ and that $\psi$ is a scalar parameter.
Notation.
[…]
2.1 Plug-in bias
We consider problems that admit regular estimators that (i) converge at the parametric $\sqrt{n}$-rate and (ii) whose asymptotic normal distributions vary smoothly with $P$. A closely related notion [53] is pathwise differentiability of the target parameter $\psi$ [20]. This means that given a smoothly parametrized one-dimensional submodel $\{P_{\epsilon}\}_{0^{-}<\epsilon<0^{+}}$ of $\mathcal{M}$, the map $\epsilon\mapsto\psi(P_{\epsilon})$ is differentiable in the ordinary sense, and the derivative has a special representation that we state below. However, before that, we need to introduce the tangent space.
Scores.
Let $h={\frac{d}{d\epsilon}}_{|\epsilon=0}\log p_{\epsilon}$ be the score of the path at $P_{0}$ (i.e., $P_{\epsilon=0}$). Then, the smoothness requirement on the path implies that $P_{0}h=0$ and $P_{0}h^{2}<\infty$ [[]Ch.7]van2000asymptotic. The score can be interpreted as the tangent vector to the path (curve) of distributions. The collection of scores of all smooth paths through $P_{0}$ forms a linear space, which we refer to as the tangent space of $\mathcal{M}$ at $P_{0}$. We will assume that the model $\mathcal{M}$ is nonparametric in the sense that for any $P$, its tangent space equals to $L^{2}_{0}(P)$. In this work, following [29], given any distribution $P\in\mathcal{M}$ and a score $h$, we will work with the concrete path
$p_{\epsilon,h}\coloneqq p(\epsilon|h)\coloneqq(1+\epsilon h)p.$ (1)
These paths are analytically convenient for us, but may not always be valid for all $\epsilon$ in an interval around 00 (e.g., if $h$ takes unbounded negative values over $\mathcal{O}$ but this will not be an issue in this work). Other constructions are available [7], but we defer these extensions to future work.
Influence function.
Returning to the differentiability of $\psi$ at $P_{0}$, it is required that there exists a continuous linear functional $D_{P_{0}}\psi:L^{2}_{0}(P_{0})\rightarrow\mathbb{R}$ such that ${\frac{d}{d\epsilon}}_{|\epsilon=0}\psi(P_{\epsilon,h})=D_{P_{0}}\psi[h]$ for every $h\in L^{2}_{0}(P_{0})$ and every smooth path $P_{\epsilon}$ with score $h$. The influence function $\tilde{\psi}_{P}\in L^{2}_{0}(P)$ of parameter $\psi$ is the Riesz representer of its derivative $D_{P}\psi$ for the $L^{2}_{0}(P)$ inner product and is characterized by the property that $D_{P}\psi[h]=\langle h\,,\,\tilde{\psi}_{P}\rangle_{L^{2}_{0}(P)}$ for every $h\in L^{2}_{0}(P)$. Intuitively, $\tilde{\psi}_{P}(o)$ measures the sensitivity of $\psi$ to perturbing $P$ toward a point-mass distribution at $o$ [37, 17]. Also note that by Cauchy-Schwarz inequality, $\tilde{\psi}_{P}$ is the direction of most rapid change in $\psi$ at $P$ [36].
The significance of $\tilde{\psi}_{P}$ for our work cannot be overstated: (i) it features in the von Mises expansion $\psi(\widehat{P})-\psi(P^{*})=-P^{*}[\tilde{\psi}_{\widehat{P}}]+R_{2}(\widehat{P},P^{*})$ which decomposes the bias of a plug-in estimator $\psi(\widehat{P})$ into the first- and second-order terms discussed in Lemma 1, (ii) it determines the asymptotic distribution [33] of regular efficient estimators of $\psi$ stated in Definition 1, and (iii) it is integral to the debiasing arguments in TMLE [29], DML [8], and this work.
Definition 1.
An estimator $\hat{\psi}$ is asymptotically linear if it satisfies $\hat{\psi}_{n}-\psi(P^{*})=\mathbb{P}_{n}\tilde{\psi}_{P^{*}}+o_{P^{*}}(1/\sqrt{n})$.
The goal of this work is to construct a distributional estimator $P^{\infty}_{n}$ such that the plug-in $\hat{\psi}_{n}=\psi(P^{\infty}_{n})$ is asymptotically linear and converges at the parametric (efficient) $\sqrt{n}$-rate to the Gaussian asymptotic distribution $N(0,P^{*}[\tilde{\psi}_{P^{*}}^{2}])$. By contrast, the estimation of the nonparametric nuisance parameter $P^{*}$ is a much harder statistical problem [51] which manifests through a strictly slower efficient rate of convergence than $\sqrt{n}$. Because of this, a naive plug-in estimator may have a first-order bias term $P^{*}[\tilde{\psi}_{\widehat{P}}]$ that dominates the $\sqrt{n}$-asymptotics. In this work, we construct $P^{\infty}_{n}$ that eliminates this bias for any $\psi$ with suitable regularity.
2.2 TMLE Preliminaries
We now briefly recap TMLE following the original paper [29] closely. The construction of our debiased estimator presented in Section 3 parallels that of TMLE.
TMLE is a plug-in estimator that satisfies Definition 1 and is constructed in two steps. First, one obtains a preliminary estimate $P^{0}_{n}$ of $P^{*}$ via, e.g., machine learning tools, kernel density estimators, etc. The plug-in estimator $\psi(P^{0}_{n})$ is then consistent but typically not efficient in the sense that it may converge at a rate slower than $\sqrt{n}$ (see e.g., examples in [29]). The TMLE solution to debiasing $P^{0}_{n}$ for the parameter $\psi$ is to perturb $P^{0}_{n}$ in a way that (i) increases the sample likelihood of the distribution estimate [^1] Recall that this is equivalent to minimizing an estimate of the KL divergence to $P^{*}$., and (ii) solves the influence curve estimating equation $\mathbb{P}_{n}\tilde{\psi}_{\widehat{P}}=0$, thereby maintaining consistency and mitigating first-order bias. This is achieved with perturbations to $P^{0}_{n}$ along the direction of the most rapid change in $\psi$. Specifically, [29] consider the parametric model in (1) with the score equal to the influence function $\tilde{\psi}_{P^{0}_{n}}$ and take the updated estimate to be the MLE $p^{1}_{n}=p^{0}_{n}(\epsilon_{n}|\tilde{\psi}_{P^{0}_{n}})$, where
$\epsilon_{n}\coloneqq\epsilon_{n}\left(\mathbb{P}_{n}|P^{0}_{n}\right)\coloneqq\operatornamewithlimits{arg\,max}_{\epsilon:p^{0}_{n}(\epsilon)\in\mathcal{M}}\sum_{i=1}^{n}\log\left(1+\epsilon\tilde{\psi}_{P^{0}_{n}}(O_{i})\right)p^{0}_{n}(O_{i}).$ (2)
Intuitively, TMLE constrains the nonparametric model $\mathcal{M}$ to a parametric submodel by disregarding all fluctuations about $p^{0}_{n}$ that do not change $\psi$ locally. With this targeted (to $\psi$) and rather extreme regularization, the TMLE step is to, very natrually, maximize the likelihood over the parametric submodel.
Assuming that the first-order conditions hold at the optimum, it follows that the updated estimate $p^{1}_{n}$ solves the following estimating equation:
$0=\mathbb{P}_{n}\frac{\frac{d}{d\epsilon}p^{0}_{n}(\epsilon)}{p^{0}_{n}(\epsilon)}=\frac{1}{n}\sum_{i=1}^{n}\frac{\tilde{\psi}_{P^{0}_{n}}(O_{i})}{1+\epsilon\tilde{\psi}_{P^{0}_{n}}(O_{i})}.$ (3)
Equation (3) is not quite the influence score equation (which it would be if $\epsilon$ were 00). However, this TMLE step can be iterated, producing a sequence of updated estimates $\{p^{k}_{n}\}_{k=0}^{\infty}$. Now, assuming that $p^{k}_{n}$ converges as $k\to\infty$ and regularity, the limit $p^{\infty}_{n}$ is a fixed point of (2), the corresponding argmax in (2) is $\epsilon_{n}(\mathbb{P}_{n}|P^{\infty}_{n})=0$, and Eq. (3) becomes the influence curve estimating equation:
$0=\mathbb{P}_{n}\tilde{\psi}_{P^{\infty}_{n}}.$ (4)
This property of $P^{\infty}_{n}$ achieves asymptotic debiasing and efficiency for the plug-in estimator $\psi(P^{\infty}_{n})$ under regularity conditions that we describe next. First, we need to slightly refine the von Mises expansion.
Lemma 1.
Let $\psi:\mathcal{M}\to\mathbb{R}$ be pathwise differentiable and $\widehat{P},P^{*}\in\mathcal{M}$. Assume that $\widehat{P}$ satisfies: (1) $P^{*}$ is absolutely continuous with respect to $\widehat{P}$, and (2) Radon–Nikodym derivative $dP^{*}/d\widehat{P}$ is square integrable under $\widehat{P}$. Consider a path $p_{\epsilon}$ with the score $h={p^{*}}/{\hat{p}}-1$ at $\epsilon=0$, e.g., $\hat{p}(\epsilon|h)=(1-\epsilon)\hat{p}+\epsilon p^{*}$. Then the following von Mises expansion holds:
$\psi(\widehat{P})-\psi(P^{*})=\mathbb{P}_{n}\tilde{\psi}_{P^{*}}-\mathbb{P}_{n}\tilde{\psi}_{\widehat{P}}+(\mathbb{P}_{n}-P^{*})\big{[}\tilde{\psi}_{\widehat{P}}-\tilde{\psi}_{P^{*}}\big{]}+R_{2}(\widehat{P},P^{*})$ (5)
where $R_{2}(\widehat{P},P^{*})$ is defined by the expansion and represents the second-order (quadratic) remainder in the difference between $\widehat{P}$ and $P^{*}$.
This expansion (proof in Appendix B) provides a decomposition of the estimation error $\psi(\widehat{P})-\psi(P^{*})$ akin to that in Definition 1: The first term $\mathbb{P}_{n}\tilde{\psi}_{P^{*}}$ governs the asymptotic distribution of $\psi(\widehat{P})$, provided that all the remaining terms are negligeable. TMLE, by construction, eliminates the plug-in bias term $\mathbb{P}_{n}\tilde{\psi}_{\widehat{P}}$. The technical assumptions we state next control the behavior of the empirical process term $(\mathbb{P}_{n}-P^{*})[\tilde{\psi}_{\widehat{P}}-\tilde{\psi}_{P^{*}}]$ and the second-order reminder term $R_{2}(\widehat{P},P^{*})$.
Assumption 1.
Assume that $\lVert\tilde{\psi}_{\widehat{P}}-\tilde{\psi}_{P^{*}}\rVert_{L^{2}(P^{*})}=o_{P^{*}}(1)$ and that there exists an event $\Omega$ with $P^{*}(\Omega)=1$ such that the class of functions $\{\tilde{\psi}_{\widehat{P}_{n}(w)}\;;\;w\in\Omega,n\in\mathbb{N}\}$ is $P^{*}$-Donsker. Then the empirical process term satisfies: $(\mathbb{P}_{n}-P^{*})\big{[}\tilde{\psi}_{\widehat{P}}-\tilde{\psi}_{P^{*}}\big{]}=o_{P^{*}}(1/\sqrt{n}).$
Assumption 2.
Assume that $\widehat{P}$ converges sufficiently fast and $\psi$ is sufficiently regular so that the second-order remainder term $R_{2}(\widehat{P},P^{*})$ is $o_{P^{*}}(1/\sqrt{n})$.
We conclude that under regularity conditions and Assumptions 1 and 2, TMLE plug-in estimator $\psi(P^{\infty}_{n})$ is asymptotically linear.
Remarks.
Works such as [28] and [29] use Assumption 1 to control the behavior of the empirical process term. To verify $L^{2}(P^{*})$ convergence, one can show that the distributional estimator $\widehat{P}$ is consistent and the influence function map $P\mapsto\tilde{\psi}_{P}$ is sufficiently smooth. Later work on TMLE [56, 9] relaxes the Donsker assumption by applying sample-splitting.
The strength Assumption 2 is dependent on the target parameter and the complexity/regularity of the model. For parameters that enjoy the widely celebrated doubly-robust property, such as the average treatment effect (ATE), this assumption is satisfied when the nuisance (e.g., the propensity score and the regression function in the case of ATE) estimators converge at an $o_{P^{*}}(n^{-1/4})$ rate [9]. For other parameters which do not have the doubly-robust property, such as the variance of the blip function, this assumption is a stronger requirement [30].
We emphasize that the influence function $\tilde{\psi}_{P}$ plays a crucial role in the theoretical analysis and in the computation of TMLE and (all, to our knowledge) other debiasing estimators. Deriving the influence function in order to implement an estimator in applied work, however, places a nontrivial burden on even the most technically trained practitioners. We propose a novel estimator $P^{\infty}_{n}$ that is debiased for any sufficiently regular parameter $\psi$ when used as a plug-in and does not require the influence function to be computed. Compared to TMLE, the proposed estimator exhibits comparable empirical performance and theoretical guarantees while requiring similar regularity assumptions.
In the construction of our estimator, we follow the TMLE framework with one main modification. Our estimator takes as input only the initial estimate $P^{0}_{n}$, which is perturbed in order to mitigate the plug-in bias. Upon convergence of the perturbation, we obtain a debiased estimate $P^{\infty}_{n}$ that eliminates the plug-in bias via mechanics similar with TMLE. Instead of reducing $\mathcal{M}$ to a one-dimensional submodels where $\psi$ changes most rapidly (which requires the influence function) in perturbing $P^{0}_{n}$, we construct RKHS-based submodels (which are infinite-dimensional and do not depend on $\tilde{\psi}_{P}$ or $\psi$). RKHSs possess two key features that facilitate our results: (i) these spaces are sufficiently rich to approximate any influence function (universality), and (ii) optimization of sample-based objectives (e.g., likelihood) over the RKHS-based models can be efficiently carried out using a representer-type theorem, allowing for tractable solutions and providing a natural regularization approach for the nonparametric MLEs. As a result, we obtain debiasing for any $\psi$ by constructing our estimator so that it satisfies the necessary estimating equation over the entire RKHS, and therefore for any influence function via universality. Thus, in order to implement our estimator, the practitioner simply needs to compute our $P^{\infty}_{n}$ and plug it into her target parameter $\psi$ without ever evaluating the influence function. Furthermore, in combination with a bootstrap scheme for inference (which we intend to explore rigorously in follow-up work), our estimation methodology completely removes the user burden of deriving the influence function in the semiparametric problems we consider.
Before we formally describe our estimator, we first review the necessary results from the RKHS literature.
2.3 RKHS Preliminaries
We start by defining a PSD kernel [^2] Not to be confused with with non-negative kernel functions, commonly used in density estimation, see [2].. Let $\mathcal{X}$ be a non-empty set. A function $K:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ is called positive-semidefinite (PSD) if for any finite sequence of inputs $\mathbf{x}=[x_{1},...,x_{n}]^{\top}\in\mathcal{X}^{n}$, the matrix $K_{\mathbf{x}}\coloneqq[K(x_{i},x_{j})]_{ij}$ is symmetric and positive-semidefinite [32].
We will need the following notation: the kernel function at $x$ is the univariate map $y\mapsto k_{x}(y)\coloneqq K(x,y)$; it is convenient and common to overload the notation with vector-valued maps $k_{\mathbf{x}}\coloneqq[k_{x_{1}},\ldots,k_{x_{n}}]^{\top}$; also, letting $\mathbf{y}=[y_{1},...,y_{m}]^{\top}$ be a column vector, we define the kernel matrix $K_{\mathbf{xy}}\coloneqq[K(x_{i},y_{j})]_{ij}\in\mathbb{R}^{n\times m}$ and $K_{\mathbf{x}}\coloneqq K_{\mathbf{xx}}$.
Next, we define a reproducing kernel Hilbert space. An RKHS is a Hilbert space $\mathcal{H}$ of real-valued functions on $\mathcal{X}$ with the property that the evaluation functional $\mathcal{H}\ni f\mapsto f(x)$ is continuous in the Hilbert norm for every $x\in\mathcal{X}$. By the Riesz representation theorem [11], every RKHS has a PSD reproducing kernel function $K:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ such that $f(x)=\langle f\,,\,k_{x}\rangle_{\mathcal{H}}$ for every $x\in\mathcal{X}$ and every $f\in\mathcal{H}$. A standard fact about RKHS is that convergence in the norm implies pointwise convergence.
Conversely, given a PSD kernel $K$, it can be shown that the space of functions $\mathrm{span}\{k_{x}\;;\;x\in\mathcal{X}\}$ with the inner product $\langle a^{\top}k_{\mathbf{x}}\,,\,b^{\top}k_{\mathbf{y}}\rangle_{K}\coloneqq a^{\top}K_{\mathbf{xy}}b,$ where $a\in\mathbb{R}^{n}$ and $b\in\mathbb{R}^{m}$, can be uniquely completed to an RKHS with reproducing kernel $K$ [55, Ch12]. This property is critical in proving our debiasing result (Theorem 2).
Property 1 (reproducing kernel Hilbert space via PSD kernel).
Let $K:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ be a PSD kernel. There exists a unique Hilbert space of functions on $\mathcal{X}$, denoted by $\mathcal{H}_{K}$, which satisfies the following properties:
  • For any $x\in\mathcal{X}$, the function $k_{x}\in\mathcal{H}_{K}$, and
  • $\langle f,k_{x}\rangle_{\mathcal{H}}=f(x)\quad\forall f\in\mathcal{H}_{K}$.
Another crucial property of an RKHS for our debised estimator is that it is a sufficiently rich space of functions in which any influence function can be suitably approximated. Let $\mathcal{C}_{0}(\mathcal{X})$ denote the space of continuous functions that vanish at infinity with the uniform norm. We say that a kernel $K$ is $\mathcal{C}_{0}$ provided that $\mathcal{H}_{K}$ is a subspace of $C_{0}(\mathcal{X})$. We say that a $\mathcal{C}_{0}$-kernel $K$ is universal if the corresponding RKHS $\mathcal{H}_{K}$ is dense in $L^{2}(\mathcal{X},P)$ for any probability measure $P$ on the sample space $\mathcal{X}$.
Property 2 (universal kernels).
Let $\mathcal{X}$ be a closed subset of $\mathbb{R}^{n}$, and let $K:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a PSD kernel such that the map $x\mapsto K(x,x)$ is bounded and $k_{x}\in\mathcal{C}_{0}(\mathcal{X})$ for every $x\in\mathcal{X}$. Then $\mathcal{H}_{K}\subset\mathcal{C}_{0}(\mathcal{X})\subset L^{p}(P)\subset L^{q}(P)$ for all $1\leq p<q\leq\infty$ and every probability measure $P$ (with continuous inclusions). The following are equivalent:
  1. 1.
    The kernel $K$ is universal;
  2. 2.
    The space $\mathcal{H}_{K}$ is dense in $\mathcal{C}_{0}(\mathcal{X})$;
  3. 3.
    The space $\mathcal{H}_{K}$ is dense in $L^{p}(\mathcal{X},P)$ for at least one (and consequently for all) $1\leq p<\infty$ and each probability measure $P$ on $\mathcal{X}$.
We refer to Carmeli et al [5] for the details.
Property 2 implies that universal kernels contain sequences that can approximate the influence function $\tilde{\psi}_{P}$ arbitrarily well with respect to the $L^{2}(P)$ norm for any $P$. One example of a universal kernel is the Gaussian kernel: $K(x,y)=\exp(-\|x-y\|_{2}^{2})$, and we will use it as the primary example throughout the rest of the paper.
Finally, our debiasing method relies on constructing local (to $P$) submodels of $\mathcal{M}$ indexed by a set of scores at $P$. Since scores have zero means, we must arrange for our RKHS to have this property. We illustrate how to construct these subspaces and their reproducing kernels in Proposition 1 (Proof in Appendix C.1).
Proposition 1 (mean-zero RKHS).
Suppose $K$ is a PSD kernel on $\mathcal{X}$ that is bounded, $C_{0}$, and universal and $P$ is a probability measure on $\mathcal{X}$.
Let $\mathcal{H}_{P}\coloneqq\{h\in\mathcal{H}_{K}\;;\;Ph=0\}$ be the subspace of $\mathcal{H}_{K}$ containing all $P$-mean-zero functions. Then $\mathcal{H}_{P}$ is closed in $\mathcal{H}_{K}$ and therefore itself is an RKHS, and
[…]
is the reproducing kernel of $\mathcal{H}_{P}$, i.e., $\mathcal{H}_{P}=\mathcal{H}_{K_{P}}$.
Furthermore, for any $h\in\mathcal{H}_{K}$, the following orthogonal decomposition with respect to $\mathcal{H}_{K}=\mathcal{H}_{P}\oplus\mathcal{H}_{P}^{\perp}$ holds:
[…]
Moreover, $\mathcal{H}_{P}$ is dense in $L^{2}_{0}(P)$.
3 Debiasing with Reproducing Kernel Hilbert Spaces
In this section, we propose a novel estimator $P_{n}^{\infty}$, which we dub as the kernel debiased plug-in estimator (KDPE), using reproducing kernel Hilbert space methods. Our estimator has three key properties: (i) it provides a plug-in estimator $\psi(P^{\infty}_{n})$ of any parameter $\psi$; (ii) similarly to TMLE, it solves the efficient influence curve estimating equation (4) asymptotically, thereby eliminating the plug-in bias of $P^{0}_{n}$ in any pathwise differentiable parameter $\psi(P)$; (iii) it does not require the computation of the influence function $\tilde{\psi}_{P}$.
First, given an initial estimate $P^{0}_{n}\in\mathcal{M}$ of $P^{*}$ and an RKHS $\mathcal{H}$, we construct the subspace $\mathcal{H}_{P^{0}_{n}}=\mathcal{H}(P^{0}_{n})=\left\{h\in\mathcal{H}\;|\;P^{0}_{n}[h]=0\right\}\subset\mathcal{T}_{P^{0}_{n}}$. By Proposition 1, $\mathcal{H}_{P^{0}_{n}}$ is itself an RKHS.
Second, we consider the collection of perturbations ~$\widetilde{\mathcal{M}}_{P_{n}^{0}}$ of the initial estimate $p^{0}_{n}$, which is constructed similarly to perturbations in TMLE (compared with (1)). We define ~$\widetilde{\mathcal{M}}_{P}$ for a distribution $P$ as follows:
[…]
Instead of considering a one-dimensional model along the influence function, we allow $p^{0}_{n}$ to be perturbed along any direction $h\in\mathcal{H}_{P^{0}_{n}}$ that produces a density in $\mathcal{M}$. This particular form of perturbation is convenient in our analysis and is well suited for $h\in\mathcal{H}$. When compared with general scores $h\in L^{2}(P)$, our RKHS scores exhibit more regularity, e.g., $h\in\mathcal{H}$ is bounded whenever the kernel is. Other RKHS-based constructions are available in the literature [13, 31, 45] and might offer computational advantages in our algorithm. We leave these extensions to future work and focus on the simplest case here for the sake of clarity.
Third, we consider the regularized empirical likelihood maximization problem with the objective function $L(p_{h},\lambda_{K})=-\sum_{i=1}^{n}\log(p_{h}(O_{i}))+\lambda_{K}\|h\|_{\mathcal{H}_{P_{h}}}^{2}$ over our RKHS-based model of perturbations to $P^{0}_{n}$:
$\operatornamewithlimits{arg\,min}_{h:p^{0}_{n,h}\in\widetilde{\mathcal{M}}_{P_{n}^{0}}}L(p_{n,h}^{0},\lambda_{K})=\operatornamewithlimits{arg\,min}_{h:p^{0}_{n,h}\in\widetilde{\mathcal{M}}_{P_{n}^{0}}}-\left[\sum_{i=1}^{n}\log(1+h(O_{i}))p^{0}_{n}(O_{i})\right]+\lambda_{K}\lVert h\rVert^{2}_{\mathcal{H}_{P_{n}^{0}}},$ (8)
where $\lambda_{K}\geq 0$ is a regularization parameter that penalizes the complexity $h$ of the perturbation to $P$. We provide theoretical results for any $\lambda_{K}\geq 0$, but in our numerical experiments, we will only consider the case where $\lambda_{K}=0$ with objective function $L(p_{h})=-\sum_{i=1}^{n}\log(1+h(O_{i}))p(O_{i})$. Stated in its most general form, the following result (proof in Appendix C.2) suggests that the general solutions of (8) are of the form $h_{n}(p^{0}_{n})(\cdot)\coloneqq\sum_{i=1}^{n}\alpha_{i,n}^{0}K_{P^{0}_{n}}(O_{i},\cdot)+\check{h}(\cdot),$ where $K_{P^{0}_{n}}$ is the kernel function of the RKHS $\mathcal{H}_{P^{0}_{n}}$.
Theorem 1 (representer).
[…]
This general solution consists of two parts that can be optimized independently of each other. The value of the objective in (8) is determined by the first component which is specified with a finite-dimensional parameter $\alpha\in\mathbb{R}^{n}$ that combines $n$ basis functions $K_{P^{0}_{n}}(O_{1},\cdot),\ldots,K_{P^{0}_{n}}(O_{n},\cdot)$; we denote by $\alpha_{n}^{0}=(\alpha_{1,n}^{0},\ldots,\alpha_{n,n}^{0})$ its optimal value. The second component $\check{h}$ can be any function in $\mathcal{H}_{P^{0}_{n}}$ such that $\check{h}(O_{i})=0$ for $i=1,\ldots,n$, so that it does not change the value of the empirical likelihood, indicating that the number of solutions to (8) is infinite. A natural way to address this issue is to choose the solution where $\check{h}\equiv 0$, making the solution unique under some regularity assumptions. [^4] Equivalently, this regularization can be viewed as the consequence of introducing an infinitesimally small penalty based on the RKHS norm, added to empirical likelihood objective in (8). The norm penalty (with any positive loading) forces $\check{h}=0$ by the RKHS representer theorem. However, in order to satisfy the efficient influence curve estimating equation (4), our method requires the norm penalty to vanish. In this work, we set the loading to 0 exactly and defer the analysis of asymptotically vanishing penalization to future work. Thus, Theorem 1 implies that to minimize (8), it suffices to consider the minimization problem over $\alpha\in\mathbb{R}^{n}$ using $\widehat{p}_{n}(\alpha,p)(\cdot)\coloneqq(1+\sum_{i=1}^{n}\alpha_{i}K_{P}(O_{i},\cdot))p(\cdot)$, an $n$-dimensional parametric model, with the corresponding loss function:
$L(\widehat{p}_{n}(\alpha,p_{n}^{0}),\lambda_{K})=-\sum_{j=1}^{n}\Big{(}\log\Big{(}\Big{(}1+\sum_{i=1}^{n}\alpha_{i}K_{P_{n}^{0}}(O_{i},O_{j})\Big{)}p_{n}^{0}(O_{j})\Big{)}\Big{)}+\lambda_{K}\Bigg{\|}\sum_{i=1}^{n}\alpha_{i}K_{P_{n}^{0}}(O_{i},\cdot)\Bigg{\|}^{2}_{\mathcal{H}_{P_{n}^{0}}}.$
Fourth, for $\lambda_{k}=0$ and $\check{h}\equiv 0$, we obtain $\alpha_{n}^{0}$ and $p_{n}^{1}$, the $1^{\text{st}}$-step of our proposed kernel debiased plug-in estimator:
$\alpha_{n}^{0}$ $=\operatornamewithlimits{arg\,min}_{\alpha:\ \widehat{p}_{n}(\alpha,p_{n}^{0})\in\widetilde{\mathcal{M}}_{P_{n}^{0}}}-\sum_{i=1}^{n}\log(\widehat{p}_{n}(\alpha,p_{n}^{0})(O_{i}),$ (9)
$p^{1}_{n}(\cdot)$ $\coloneqq\widehat{p}_{n}(\alpha_{n}^{0},p_{n}^{0})=\Big{(}1+\sum_{i=1}^{n}\alpha_{i,n}^{0}K_{P^{0}_{n}}(O_{i},\cdot)\Big{)}p^{0}_{n}(\cdot),$ (10)
and the $k+1^{\text{st}}$-step of KDPE by $p^{k+1}_{n}\coloneqq\widehat{p}_{n}(\alpha_{n}^{k},p_{n}^{k})$, recursively. KDPE terminates when $\alpha_{n}^{k}=0$, which indicates $\widehat{p}_{n}(0,p_{n}^{k})$ is a local optimum for $L(\cdot)$. We provide the pseudocode for KDPE in Algorithm 1.
Algorithm 1 Pseudo-Code for KDPE ($\lambda_{K}=0$)
Initialize the starting distribution ${P}^{0}_{n}$, $\alpha=\infty$, and $i=0$.
while $\alpha_{n}^{i}\neq 0$ do
     Update the distribution according to the objective function $L(\cdot)$:
[…]
     ${p}_{n}^{i+1}=\widehat{p}_{n}(\alpha^{i}_{n},{p}^{i}_{n})$, and $i=i+1$
end while
Return $\psi({P}_{n}^{k})$, where $k$ is the number of iterations needed until termination.
Similar to TMLE, we make the following assumption about the asymptotic behavior with respect to $k$:
Assumption 3 (convergence and finite termination at a interior point).
We assume that the $k$-step KDPE estimate $p_{n}^{k}$ converges almost surely to some limit $p_{n}^{\infty}$ in the interior of $\mathcal{M}$, and that this limiting distribution $P_{n}^{\infty}\equiv\widehat{P}$ is a local minimum of (8) in the interior of the perturbation space ~$\widetilde{\mathcal{M}}_{\widehat{P}}$.
In Assumption 3, we assume that our procedure converges to a limit $P_{n}^{\infty}$, and that this limit $P_{n}^{\infty}$ satisfies two interior conditions: 1) the interior condition on the full model class $\mathcal{M}$ to ensure that our problem is well-defined, [^5] That is, our problem is nonparametric, i.e., the tangent space $\mathcal{T}_{P_{n}^{\infty}}=L_{0}^{2}(P_{n}^{\infty})$, and our target parameter is pathwise differentiable, i.e., the conditions in Lemma 1 are satisfied. and 2) the interior condition on the submodel ~$\widetilde{\mathcal{M}}_{P_{n}^{\infty}}$ in order for the first-order optimality conditions hold. The latter is needed for the following debiasing result (Proposition 2) to hold. We stress that this assumption is commonly adopted in TMLE literature [29] (see a more detailed discussion in Remark C.1). Under Assumption 3, Proposition 2 (proof in Appendix C.3) provides a key consequence of KDPE:
Proposition 2.
Under Assumption 3, the final KDPE estimate $P_{n}^{\infty}=\widehat{P}$ of the true distribution $P^{*}$ satisfies the following estimating equations:
$\mathbb{P}_{n}h=0\quad\text{for all }h\in\mathcal{H}_{\widehat{P}}.$ (11)
This result is somewhat striking at first. It is helpful to consider it in duality with the solution (9) to the optimization problem in (8). In (8) we solve $n$ first-order conditions with an infinite-dimensional parameter $h\in\mathcal{H}_{P}$ with $P$ fixed, which results in a multiplicity of solutions. By contrast, (9) fixes an infinite dimensional ‘free variable’ $\check{h}=0$, and reduces the problem to $n$ equations in $n$ unknowns. By the first-order gradient condition guaranteed by Assumption 3, KDPE solves the infinite system of equations posed in Proposition 2 at the limiting distribution $\widehat{P}$ by iteratively performing the update step presented in (9).
3.1 Universal Kernels for KDPE
While Proposition 2 states that KDPE solves infinitely many score equations, it only guarantees that the plug-in bias, $\mathbb{P}_{n}\tilde{\psi}_{\widehat{P}}$, is $o_{P^{*}}(1/\sqrt{n})$ if $\tilde{\psi}_{\widehat{P}}\in\mathcal{H}_{\widehat{P}}$ $P^{*}$-almost surely. By choosing a universal kernel [5, 32, 57], we strengthen the results of Proposition 2 to achieve our desired result, under the regularity conditions below.
Assumption 4.
We assume that (1) the estimate $\widehat{P}$ satisfies $p^{*}/\widehat{p}\in L^{2}(P^{*})$ with the estimated influence function satisfying $\tilde{\psi}_{\widehat{P}}\in L^{2}(P^{*})$ $P^{*}$-almost surely, and that (2) the kernel function $K$ is universal.
The first condition in Assumption 4 is a common assumption for debiasing methods [29], and serves two technical purposes for KDPE. It (1) enables a change of measure from $\widehat{P}$ to $P^{*}$ that maintains square integrability, and (2) guarantees the conditions necessary for the expansion in Lemma 1. Because $\widehat{p}/p^{*}$ is square integrable, this also implies that $\tilde{\psi}(P^{*})\in L_{0}^{2}(\widehat{P})\cap L^{2}(P^{*})$. The second condition in this assumption is satisfied by using a universal kernel, such as the mean-zero Gaussian kernel that satisfies the conditions in Proposition 1. Combining both parts of Assumption 4 implies that $\mathcal{H}_{\widehat{P}}$ is dense in the set of functions $L_{0}^{2}(\widehat{P})\cap L^{2}(P^{*})$ under the norm $\|\cdot\|_{L_{2}(P^{*})}$: this crucially guarantees the existence of an approximating sequence $\{h_{j}\}_{j\in\mathbb{N}}\subset\mathcal{H}_{K}$ that is arbitrarily close to $\tilde{\psi}_{\widehat{P}}$ with respect to the ${L_{2}(P^{*})}$ norm. Lastly, we provide a sufficient condition for KDPE, which closely resembles the Donsker assumptions in Assumption 1:
Assumption 5.
Assume there exist: (i) an event $\Omega$ with $P^{*}(\Omega)=1$, and, for every $\omega\in\Omega$, (ii) a sequence $h_{j}(\omega)\in\mathcal{H}_{\widehat{P}(\omega)}$ indexed by $\omega\in\Omega$ such that $\lVert h_{j}(\omega)-\tilde{\psi}_{\widehat{P}(\omega)}\rVert_{L^{2}(P^{*})}\xrightarrow[j\to\infty]{}0$ uniformly in $w\in\Omega$ and the set of functions $\{h_{j}(\omega)\;;\;\omega\in\Omega,j\in\mathbb{N}\}$ is $P^{*}$-Donsker as $n,j\rightarrow\infty$.
Assumption 5 is our main regularity assumption on the model $\mathcal{M}$ and functional $\psi$ needed to control the plug-in bias term $\mathbb{P}_{n}\tilde{\psi}_{\widehat{P}}$ of the proposed KDPE estimator. It requires that the set of influence functions =$\mathcal{F}=\{\tilde{\psi}_{\widehat{P}}\}_{\widehat{P}}$ be sufficiently small/regular and that this property can be projected to those scores in $\mathcal{H}_{K}$ that are sufficiently close to $\mathcal{F}$  [^6] While this assumption provides a clear intuition for the debiasing mechanism of KDPE, it is not strictly required for the method to work. Less extreme conditions on the regularity of $\mathcal{F}$, such as $C^{\alpha}$ or Sobolev norm bounds (see van der Vaart [52, Ch 19]) and RKHS approximation theory [49]), can imply the Donsker property in Assumption 5.. Next, we formally state our main result (proof in Appendix C.4):
Theorem 2 (asymptotic linearity of KDPE).
Let $\psi:\mathcal{M}\to\mathbb{R}$ be a smooth functional of distribution $P$ with influence function $\tilde{\psi}_{P}\in L^{2}_{0}(P)$ and von Mises expansion:
$\psi(\bar{P})-\psi(P)=\int\tilde{\psi}_{\bar{P}}d(\bar{P}-P)+R_{2}(\bar{P},P)\quad\text{for any }\bar{P},P\in\mathcal{M},$
which defines the second-order reminder term $R_{2}(\bar{P},P)$. Then, under Assumptions 1, 2,3, 4, 5,
$\psi(\widehat{P})-\psi(P^{*})=\mathbb{P}_{n}\tilde{\psi}_{P^{*}}+o_{P^{*}}(1/\sqrt{n}).$
By combining all of the necessary regularity conditions, Theorem 2 concludes that $\psi(P^{\infty}_{n})$ is asymptotically linear. We include further discussion of the assumptions in Remark C.2. In Appendix LABEL:app:ate_theory, we provide an illustration of Theorem 2 for the special case of the target parameter ATE.
4 Practical Algorithm and Simulation Results
In Section 3, we provide the theoretical foundations for debiasing via universal kernels. In this section, we provide a modified algorithm, Algorithm 2, that includes an early stopping criterion to address computational concerns. Within the context of a simulated data-generating process, we demonstrate empirical results that (i) provide guidelines for choosing hyperparameters, (ii) validate the theoretical results provided by Theorem 2, and (iii) show a simple bootstrap scheme for EIF-free inference.
4.1 Simulation Set-Up
The setup of this simulation largely follows classical examples of TMLE, including the update mechanism. For more details regarding this set-up, we refer readers to [27, Ch.4].
Data-Generating Process and Target Estimands.
We define $\mathcal{O}\equiv\mathcal{X}\times\mathcal{A}\times\mathcal{Y}$, with baseline covariate $X\in\mathcal{X}\equiv[0,1]$, binary treatment indicator $A\in\mathcal{A}\equiv\{0,1\}$, and a binary outcome variable $Y\in\mathcal{Y}\equiv\{0,1\}$. The true distribution $P^{*}=P^{*}_{X}\times P^{*}_{A|X}\times P^{*}_{Y|A,X})$ and is specified as follows: $P^{*}(X)\sim\text{Unif}[0,1]$, $P^{*}(A|X)\sim\text{Bern}\left(0.5+\frac{1}{3}\sin\left({50X}/{\pi}\right)\right)$, $P^{*}(Y|A,X)\sim\text{Bern}\left(0.4+A(X-0.3)^{2}+\frac{1}{4}\sin({40X}/{\pi})\right)$.
To impose a causal interpretation of our estimates, we assume the standard conditions of consistency, exchangability, positivity, etc. Denoting $\mu_{1}(P)=P_{X}[P_{Y|A,X}[Y|A=1,X=x]]$, and $\mu_{0}(P)=P_{X}[P_{Y|A,X}[Y|A=0,X=x]]$, the target parameters that we will investigate are 1) relative risk, $\psi_{\textsf{RR}}(P)=\mu_{1}(P)/\mu_{0}(P)$, 2) odds ratio, $\psi_{\textsf{OR}}(P)=\frac{\mu_{1}(P)/(1-\mu_{1}(P))}{\mu_{0}(P)/(1-\mu_{0}(P))}$, and 3) average treatment effect, $\psi_{\textsf{ATE}}(P)=\mu_{1}(P)-\mu_{0}(P).$
Baseline Methods for Comparison.
We compute these estimators with three different plug-in methods: 1) our proposed method KDPE, 2) TMLE, and 3) a baseline using the biased super-learned distribution, SL. A detailed description of the baseline methods is included in Appendix D.1. Comparisons against SL [41] illustrate the effect of the KDPE approach on the distribution of the estimator. TMLE, known for its asymptotic normality with the correct limiting variance, serves as a benchmark for evaluating the performance of KDPE. For TMLE, we learn a distinct data-generating distribution for each target parameter, i.e., $\widehat{P}_{\text{TMLE}}^{\mathrm{RR}},\widehat{P}_{\text{TMLE}}^{\mathrm{OR}},\widehat{P}_{\text{TMLE}}^{\mathrm{ATE}}$, to debias the corresponding estimator. For KDPE and SL, each only uses a single estimated distribution, and we denote them $\widehat{P}_{\text{SL}}$ and $\widehat{P}_{\text{KDPE}}$, respectively.
Update Step for KDPE.
We follow the classical perturbation set-up for TMLE, which begins with a superlearned [24] distribution ${P}_{n}^{0}$. For ${P}_{n}^{0}$, the initial distribution estimates the propensity score function ${P}_{n}^{0}(A|X)$ and the conditional regression function ${P}_{n}^{0}(Y|A,X)$, and sets the marginal distribution ${P}_{n}^{0}(X)=\mathbb{P}_{n}(X)$, where $\mathbb{P}_{n}(X)$ is the empirical distribution of $X$. As is standard for TMLE, we only perturb the conditional density function $p_{Y|A,X}$, using the $n$-dimensional parametric submodel $\widehat{p}_{n}(\alpha,p_{n}^{i})=(1+\sum_{i=1}^{n}\alpha_{i}K(O_{i},\cdot)){p}_{n}^{i}(Y|A,X)$.
Modifications for KDPE.
To make use of practical solvers, we work with a slightly modified algorithm, Algorithm 2. The main modifications from the baseline KDPE include the incorporation of a density bound parameter, $c$, and a convergence tolerance function. The constant $c$ preserves the condition of absolute continuity of $P^{*}$ with respect to the final distribution estimate and enables standard convex/concave optimization software for finding $\alpha^{i+1}$. The convergence tolerance function, $l(\cdot,\cdot)$ acts as a metric tracking the net change in distribution between iterations. Because the objective function $L(\cdot)=-n\times\mathbb{P}_{n}\log(\cdot)$ is globally convex, when $l(P_{n}^{i+1},P_{n}^{i})$ is small, we are close to a solution satisfying the necessary first-order optimality criteria. Examples of convergence tolerance functions include $l(P,P^{\prime})=||P-P^{\prime}||_{L_{2}(\lambda)}^{2}$, where $\lambda$ is a dominating measure, $||\alpha||_{2}^{2}$, or other distributional distance metrics.
Simulation Settings.
In all experiments, we use the number of samples $n=300$. For our modified KDPE algorithm, we use the mean-zero Gaussian kernel, [^7] The mean-zero Gaussian kernel is dense with respect to $L_{0}^{2}(P)$ for any probability measure $P$ on $\mathcal{O}$ under the conditions of Proposition 1, and is a natural choice to demonstrate the theoretical results in Section 3. We note that by fixing $P_{n}^{0}(X)=\mathbb{P}_{n}(X)$ (as is standard in TMLE), we avoid the issue of numerically approximating integrals, which naturally arise in constructing mean zero kernels. the model $\widehat{p}_{n}(\alpha,p)=(1+\sum_{i=1}^{n}\alpha_{i}K_{P}(O_{i},\cdot))p_{Y|A,X}(\cdot)$, where $p_{Y|A,X}$ denotes the conditional density of $P$ w.r.t. the counting measure $\lambda$, the log-likelihood loss function $L(P)=-n\times\mathbb{P}_{n}\log(p)$, and the convergence tolerance function $l(P,P^{\prime})=\|P-P^{\prime}\|_{L_{2}(\lambda)}^{2}$. For evaluating the effect of the hyperparameters in Section 4.2, we use 100 simulations. To evaluate the distribution of our KDPE estimator in Section 4.3, we use 450 simulations, with parameters $c=0.002$, $\gamma=0.001$.
Algorithm 2 Pseudo-Code for Modified KDPE
Input: $\{o_{i}\}_{i=1}^{n}$, convergence tolerance function $l(\cdot)$, convergence tolerance $\gamma$, density bound $c$.
Using the SuperLearner algorithm $S(\cdot)$, set ${P}^{0}_{n}=S(\{O_{i}\}_{i=1}^{n})$.
Set the convergence tracker $\beta(P^{0}_{n})=\infty$ and $i=0$.
while $\beta(P_{n}^{i})>\gamma$ do
     Update the distribution according to the loss function $L(\cdot)$:
     Set $\alpha^{i+1}=\operatornamewithlimits{arg\,min}_{\alpha\in\mathbb{R}^{n},\ c\leq\widehat{p}_{n}(\alpha,{p}_{n}^{i})\leq 1-c}L(\widehat{p}_{n}(\alpha,p_{n}^{i}))$
     Set ${p}_{n}^{i+1}=\widehat{p}_{n}(\alpha^{i+1},{p}^{i}_{n})$
     Set $\beta({P}_{n}^{i+1})=l(P_{n}^{i+1},{P}_{n}^{i})$, and $i=i+1$.
end while
Return $\widehat{P}_{\text{KDPE}}={P}_{n}^{i}$.
4.2 Selection of Hyperparameter Values
Figure 1: Density histogram of $\psi_{\mathrm{ATE}}(\widehat{P}_{\text{KDPE}})$ as a function of $c,\gamma$. $X$-axis: ATE values. $Y$-axis: density. Blue dotted line: the true ATE under $P^{*}$.
The practical modifications for KDPE pose two potential concerns regarding the conditions necessary for Theorem 2. Large values of the density bound $c$ may result in the algorithm terminating at the boundary of the perturbation space, violating the interior optimality condition. Large values of the convergence tolerance $\gamma$ may result in premature termination, resulting in a final estimate $\widehat{P}_{\text{KDPE}}$ far from an optimal solution. All theoretical results in Section 3 are derived under the setting of $c=0,\gamma=0$, and thus suggest that these parameters should be set as small as computationally feasible.
To empirically evaluate the effect of $c,\gamma$ on the distribution of our estimate, we test a grid of values $(c,\gamma)=\{0.05,0.01,0.002\}\times\{0.025,0.005,0.001\}$, and plot the simulated distribution of $\psi_{\mathrm{ATE}}(\widehat{P}_{\text{KDPE}})$ for each pair of hyperparameters in Figure 1. These empirical results confirm that the $c,\gamma$ should both be set as small as possible. The pair $(c=0.002,\gamma=0.001)$ achieves the closest simulated distribution to the asymptotically normal limiting distribution. For large values of the density bound ($c=0.05$) or convergence tolerance ($\gamma=0.025$), the simulated distributions of $\psi_{\mathrm{ATE}}(\widehat{P}_{\text{KDPE}})$ are skewed and/or possess multiple modes, showing violations of the necessary regularity conditions of Theorem 2. We discuss the choice of $c$ and $\gamma$ further in Appendix D.2.
From a practical perspective, $\gamma$ contributes most to the computational cost of Algorithm 2. For all tested values of $c$, halving the convergence tolerance $\gamma$ roughly corresponds to doubling the number of average iterations; by contrast, halving the density bound $c$ minimally changes the average number of iterations. Thus, in practice, we recommend setting $c$ as small as possible, and setting $\gamma$ as small as computationally feasible.
4.3 Empirical Distribution of Plug-in Estimators
Figure 2: Simulated distributions of $\widehat{\psi}_{\mathrm{ATE}}$ compared to their asymptotic distributions.
Figure 3: Simulated distributions of $\widehat{\psi}_{\mathrm{OR}},\widehat{\psi}_{\mathrm{RR}},\widehat{\psi}_{\mathrm{ATE}}$ for KDPE, SL, and TMLE.
Using the minimum values for the parameters tested in Section 4.2 ($c=0.002,\gamma=0.001$), we verify the theoretical guarantees of Section 3. In paticular, our simulations demonstrate two distinct results of Theorem 2: (1) KDPE achieves asymptotic normality with the minimum limiting distribution, and (2) KDPE works well as a plug-in distribution for all pathwise differentiable parameters.
As an example, the limiting distribution of $\psi_{\mathrm{ATE}}(P^{*})$ is given by $N(\psi_{\mathrm{ATE}}(P^{*}),P^{*}[\tilde{\psi}_{\mathrm{ATE}}(P^{*})^{2}])\approx N(0.1233,$ 0.0025)0.0025)0.0025 ) in our setup. Figure 2 plots the simulated distributions for $\psi_{\mathrm{ATE}}(\widehat{P}_{\text{KDPE}}),$ $\psi_{\mathrm{ATE}}(\widehat{P}_{\text{TMLE}}^{\mathrm{ATE}}),$ $\psi_{\mathrm{ATE}}(\widehat{P}_{\text{SL}})$ against the density of this limiting distribution. The distribution of $\psi_{\mathrm{ATE}}(\widehat{P}_{\text{SL}})$ in Figure 2 (and $\psi_{\mathrm{RR}}(\widehat{P}_{\mathrm{SL}}),\psi_{\mathrm{OR}}(\widehat{P}_{\text{SL}})$ in Figure 3) highlight the necessity of debiasing. Without correction, the estimated distributions deviate significantly from the asymptotic limiting distribution. In contrast, both $\psi_{\mathrm{ATE}}(\widehat{P}_{\text{KDPE}})$ and $\psi_{\mathrm{ATE}}(\widehat{P}_{\text{TMLE}}^{\mathrm{ATE}})$ achieve similar distributions to the limiting distribution. Despite no knowledge of the influence function when performing debiasing, the simulated distribution of $\psi_{\mathrm{ATE}}(\widehat{P}_{\text{KDPE}})$ closely resembles the efficient limiting distribution, corroborating the results of Theorem 2.
Figure 3 verifies the second consequence of Theorem 2: for all pathwise differentiable parameters, $\widehat{P}_{\text{KDPE}}$ is debiased as a plug-in estimator. In particular, we see that across all three target estimands, the distributions of the plug-in estimates with $\psi(\widehat{P}_{\text{KDPE}})$ perform almost identically compared with TMLE, which uses a different targeted plug-in distribution ($\widehat{P}_{\text{TMLE}}^{\mathrm{ATE}},\widehat{P}_{\text{TMLE}}^{\mathrm{RR}},\widehat{P}_{\text{TMLE}}^{\mathrm{OR}}$, respectively) for each estimand.
4.4 Bootstrap Intervals for Inference
To conduct inference, we use the classical bootstrap procedure [21], i.e., sampling $n$ observations from our observed data $\{o_{i}\}_{i=1}^{n}$ with replacement, to estimate the variance of our estimator, and use this as a plug-in estimate to construct normal approximations for the confidence interval. A detailed description of the bootstrap procedure is included in Appendix D.3. The bootstrap is a natural choice for our approach, as it does not rely on knowledge of the influence function to approximate this variance. As a proof of concept, Table 1 demonstrates empirical results demonstrating the KDPE bootstrap-estimated variance using 250 bootstrap simulations with $c=0.002,\gamma=0.001$, on a single draw of $n=300$ samples. For comparison, we include the TMLE-estimated variance, which estimates the variance $\widehat{S}(\widehat{\psi})=\mathbb{P}_{n}(\tilde{\psi}_{\widehat{P}_{\text{TMLE}}}^{\text{target}})$ using knowledge of the EIF. The baselines for both TMLE and KDPE are the sample variances of the simulated distributions from Section 4.3.
Method $\widehat{S}(\widehat{\psi}_{\text{ATE}})$ $\widehat{S}(\widehat{\psi}_{\text{RR}})$ $\widehat{S}(\widehat{\psi}_{\text{OR}})$
KDPE Bootstrap 0.00441 0.05365 0.36737
Baseline 0.00287 0.02555 0.15322
TMLE IF-Based 0.00220 0.01383 0.06105
Baseline 0.00296 0.02837 0.14981
Table 1: Estimated variance of $\widehat{\psi}_{\text{ATE}}$, $\widehat{\psi}_{\text{RR}}$, $\widehat{\psi}_{\text{RR}}$ of KDPE and TMLE on one simulation.
While the baseline variances for KDPE and TMLE are remarkably similar, the estimated variances with KDPE and TMLE differ significantly. The KDPE-bootstrap variance estimates overestimate the variance for all target parameters relative to the KDPE baseline; by contrast, TMLE’s IF-based variance estimate underestimates the variance for all target parameters. Thus, using the bootstrap approach may lead to conservative confidence intervals, and suggest that future work is necessary in order to calibrate the bootstrap confidence intervals. We stress that these results follow from a single draw of $n=300$ from $P^{*}$, so these trends are by no means conclusive. Rather, they provide a natural starting point for future work. We include additional computational results on bootstrap confidence intervals in Appendix D.4.
5 Discussion, Limitations, and Conclusions
Future work includes (i) verifying the validity of bootstrap, (ii) analyzing the effect of RKHS-norm regularization on the convergence and efficiency, (iii) investigating how to apply sample-splitting and the extent to which it can relax our assumptions, (iv) providing primitive conditions for RKHS approximation of the efficient influence estimating equation, (v) extensions into finite-sample-valid confidence intervals (as used for HAL-TMLE procedures). Alternative methods for constructing confidence intervals include directly taking the quantiles of the bootstrap estimates and/or subsampling while bootstrapping, which we do not explore in this project. The strong connection between our method and TMLE raises the question of whether variants of TMLE, such as CTMLE [9], CV-TMLE [56], and one-step TMLE [28], can be adapted to our approach. CV-TMLE is of particular interest, as it could relax the Donsker conditions in Assumptions 1 and 5. Exploring different loss functions, such as MSE, as utilized in TMLE, could provide analytical solutions for the optimiuzation step, eliminating the need for solvers and mitigating numerical instability concerns.
Acknowledgement
The authors would like to thank Susan Murphy for her early feedback and Jan-Christian Hütter for stimulating and insightful discussion. Research reported in this manuscript is supported by NIH/NIDA P50 DA054039, NIH/NIBIB and OD P41 EB028242.
Appendix A Notation
$\mathcal{O}$
Sample space, with $\mathcal{O}\subset\mathbb{R}^{d}$ and Borel $\sigma$-algebra $\mathcal{B}$;
$\mathcal{M}$
Statistical model, collection of probability measures on $(\mathcal{O},\mathcal{B})$;
$\psi(P)$
Target parameter functional, mapping from $\mathcal{M}\rightarrow\mathbb{R}$
$\tilde{\psi}_{P}$
Influence function, mapping from $\mathcal{O}\rightarrow\mathbb{R}$
$Pf,P{[}f{]}$
[…]
$\mathbb{P}_{n},\mathbb{P}_{n}f$
Empirical distribution function
$L^{2}_{0}(P)$
The set of measurable functions $\{f:\mathcal{O}\rightarrow\mathbb{R}:Pf=0\ ,Pf^{2}<\infty\}$
$P^{*}$
(Unknown) data generating distribution;
$K,\mathcal{H}_{K}$
The PSD kernel and associated RKHS with kernel function $K(\cdot,\cdot)$
$k_{x}$
The function $k_{x}\coloneqq K(x,\cdot):\mathcal{X}\rightarrow\mathbb{R}$, for $x$ in support $\mathcal{X}$
$k_{\mathbf{x}}$
Vectorized version of function $k_{x}$, where $k_{\mathbf{x}}\coloneqq[k_{x_{1}},...,k_{x_{n}}]^{T}$
$K_{\mathbf{xy}},K_{\mathbf{x}}$
Matrix of values $K_{\mathbf{x}}\coloneqq[K(x_{i},x_{j})]_{i,j}\in\mathbb{R}{n\times n}$, $K_{\mathbf{xy}}\coloneqq[K(x_{i},y_{j})]_{ij}\in\mathbb{R}{n\times m}$
$K_{P},\mathcal{H}_{P}$
The mean-zero kernel (with respect to $P$) and associated RKHS
~$\widetilde{\mathcal{M}}_{P}$
The collection of perturbations ~$\widetilde{\mathcal{M}}_{P}\coloneqq\{(1+h)p\ |\ h\in\mathcal{H}_{P}\text{ such that }(1+h)p>0\}\cap\mathcal{M}$
$p(\epsilon|h),\ p_{\epsilon,h}$
1-dimensional parametric model $(1+\epsilon h)p$
$D_{P}\psi$
Derivative functional (assumed to be continuous, linear) that maps $L_{0}^{2}(P)\rightarrow\mathbb{R}$,
$\frac{d}{d\epsilon}\psi({P_{\epsilon,h}})|_{\epsilon=0}=D_{P}\psi[h]$ for all $h\in L_{0}^{2}(P)$
$p_{h}$
Infinite-dimensional RKHS model $(1+h)p$, where $h\in\mathcal{H}_{P}$
$L(p_{h},\lambda_{K})$
Regularized log-likelihood objective function $-\sum_{i=1}^{n}\log(p_{h}(O_{i}))+\lambda_{K}\|h\|_{\mathcal{H}_{P_{h}}}^{2}$
$\widehat{p}_{n}(\alpha,p)$
$n$-dimensional parametric model $(1+\sum_{i=1}^{n}\alpha_{i}K_{P}(O_{i},\cdot))p(\cdot)$
$L(\widehat{p}_{n}(\alpha,p),\lambda_{K})$
Regularized log-likelihood objective function with $n$-dimensional parametric model:
$L(\widehat{p}_{n}(\alpha,p),\lambda_{K})=-\sum_{j=1}^{n}\Big{(}\log\Big{(}\Big{(}1+\sum_{i=1}^{n}\alpha_{i}K_{P}(O_{i},O_{j})\Big{)}p(O_{j})\Big{)}\Big{)}+\lambda_{K}\|\sum_{i=1}^{n}\alpha_{i}K_{P}(O_{i},\cdot)\|^{2}_{\mathcal{H}_{P}}$
$L(\widehat{p}_{n}(\alpha,p))$
Log-likelihood objective function $L(\widehat{p}_{n}(\alpha,p))\coloneqq L(\widehat{p}_{n}(\alpha,p),0)$
$\pi_{S}$
The orthogonal projection operator $\mathcal{H}\rightarrow S$, projects a function $h\in\mathcal{H}$ onto a set of functions $S$
$P_{n}^{i},\ p_{n}^{i}$
Estimated data distribution of TMLE/KDPE at iteration $i$
$\alpha_{n}^{i}$
The optimal perturbation for objective function $L(\widehat{p}_{n}(\alpha,p),\lambda_{K})$ at iteration $i$ for KDPE
$P^{\infty}_{n},\ p_{n}^{\infty}$
Limiting/final estimate for TMLE/KDPE
$\widehat{P}$
Generic estimated distribution, subscripts/superscript added accordingly
$\widehat{\psi}_{\text{param.}}$
Estimated target parameter value $\psi_{\text{param.}}(\widehat{P})$ using plug-in distribution $\widehat{P}$, subscripts/superscript added accordingly
$Q(P)$
Relevant portion of distribution (e.g. $\{Q_{a}(X)\}_{a\in\{0,1\}}=\{P[Y=1|X,A=a]\}_{a\in\{0,1\}}$).
$\Omega$
An event (or set of events) such that has probability 1 under $P^{*}$, i.e. $P^{*}(\Omega)=1$. The standard example is the set of all possible $n$-sample realizations $\{\{O_{i}(w)\}_{i=1}^{n}\ :\ w\in\Omega\}$ from distribution $P^{*}$.
$f(w),P(w)$
The random function / distribution, indexed by events $w\in\Omega$. Often, we use $w\in\Omega$ to index an $n$-sample realization, $\{O_{i}(w)\}_{i=1}^{n}$, from DGP $P^{*}$.
Appendix B Proof of Lemma 1
This expansion plays a significant role in motivating current estimation methods, such as [29] and [4], and it also plays a crucial role in our method as well. We provide the proof for this expansion below:
Proof of Lemma 1.
We begin by revisiting a few simple results that enable this form of expansion:
  1. L.1
    By defining $h=p^{*}/\widehat{p}-1$, we get that $h$ is mean zero under $\widehat{P}$:
[…]
    Furthermore, if $f$ is an integrable function with respect to measure $P^{*}$, then:
[…]
  2. L.2
    Since $P^{*}$ is absolutely continuous with respect to $\widehat{P}$, we have that $P_{\epsilon=1,h}=P^{*}$:
    $p(\epsilon=1|h)=(1+p^{*}/\widehat{p}-1)\widehat{p}=p^{*}$
  3. L.3
    Recall by the definition of $\mathcal{T}_{{P}}$, ${P}[\tilde{\psi}_{{P}}]=0$ for any $P$ in the model.
Consider the first-order distributional Taylor expansion of pathwise differentiable $\psi$ along $\hat{p}_{\epsilon,h}$ around $\hat{p}$ (i.e., $\epsilon=0$); note that the remainder term is defined by this expansion:
$-[\psi(\widehat{P})-\psi(P^{*})]$ $=\frac{\psi(P_{\epsilon,h})-\psi(\widehat{P})}{\epsilon}\big{|}_{\epsilon=1}$
$\eqqcolon D_{\widehat{P}}\psi[h]+R_{2}(P^{*},\widehat{P})$
$=\widehat{P}[\tilde{\psi}_{\widehat{P}}h]+R_{2}(P^{*},\widehat{P})$
$=\widehat{P}[\tilde{\psi}_{\widehat{P}}h]-\widehat{P}[\tilde{\psi}_{\widehat{P}}]+R_{2}(P^{*},\widehat{P})$
$=\widehat{P}[\tilde{\psi}_{\widehat{P}}(1+h)]+R_{2}(P^{*},\widehat{P})$
$=P^{*}[\tilde{\psi}_{\widehat{P}}]+R_{2}(P^{*},\widehat{P})$ (12)
Then, we subtract the following zero term in our expansion
$\mathbb{P}_{n}(\tilde{\psi}_{P^{*}})-\mathbb{P}_{n}(\tilde{\psi}_{\widehat{P}})+(\mathbb{P}_{n}-P^{*})(-\tilde{\psi}_{P^{*}})+\mathbb{P}_{n}(\tilde{\psi}_{\widehat{P}}).$
Collecting terms, multiplying by $-1$ and setting $R_{2}(\widehat{P},P^{*})=-R_{2}(P^{*},\widehat{P})$ leads to:
$\psi(\widehat{P})-\psi(P^{*})$
$\qquad=-P^{*}[\tilde{\psi}_{\widehat{P}}]+R_{2}(\widehat{P},P^{*})+\Big{[}\mathbb{P}_{n}(\tilde{\psi}_{P^{*}})-\mathbb{P}_{n}(\tilde{\psi}_{\widehat{P}})+(\mathbb{P}_{n}-P^{*})(-\tilde{\psi}_{P^{*}})+\mathbb{P}_{n}(\tilde{\psi}_{\widehat{P}})\Big{]}$
$\qquad=\mathbb{P}_{n}\tilde{\psi}_{P^{*}}-\mathbb{P}_{n}\tilde{\psi}_{\widehat{P}}+(\mathbb{P}_{n}-P^{*})\big{[}\tilde{\psi}_{\widehat{P}}-\tilde{\psi}_{P^{*}}\big{]}+R_{2}(\widehat{P},P^{*}).$ (5)
Appendix C KDPE Theory
C.1 Proof of Proposition 1
Proof of Proposition 1.
Because $K$ is bounded, the expectation functional $h\mapsto\mathbb{E}_{P}[h]$ is linear and continuous on $\mathcal{H}_{K}$. Next, we note that $\varphi\in\mathcal{H}_{K}$ and it is the Riesz representer of the expectation functional (we have shown this elsewhere and remark here that the rigorous argument considers a Riemann approximation to the expectation integral and weakly convergent subsequences). It follows that $h\in\mathcal{H}_{P}$ if and only if $\langle h\,,\,\varphi\rangle_{K}=0$, and from this we can deduce the projection operator onto $\mathcal{H}_{P}$ and the reproducing kernel of that subspace.
Finally, given $h\in L^{2}_{0}(P)$, let $\{g_{n}\}\subset\mathcal{H}_{K}$ be an approximating sequence of $h$ in $\mathcal{H}_{k}$. Let $g_{n}\tilde{g}_{n}+\mathbb{E}_{P}[g_{n}]/\mathbb{E}_{P\times P}[K]\varphi$ be the RKHS orthogonal decomposition of $g_{n}$ into $\mathcal{H}_{P}$ and $\mathcal{H}_{P}^{\perp}$, and note that this operation is continuous on $L^{2}(P)$. By continuity of expectation in $L^{2}(P)$, $\mathbb{E}_{P}[g_{n}]\to 0$ in $L^{2}(P)$. It follows that the orthogonal part of $g_{n}$ vanishes in $L^{2}(P)$ and $\tilde{g}_{n}\to h$ in $L^{2}(P)$. Thus, $\mathcal{H}_{P}$ is dense in $L^{2}_{0}(P)$. ∎
C.2 Proof of Theorem 1
Let $L(Q(P)):\mathcal{X}\to\mathbb{R}$ be a general loss function that identifies the relevant part, $Q(P)$, for the target parameter of interest $\psi(P)=\psi(Q(P))$, satisfying the property that $\mathbb{E}_{P}L(Q(P))\leq\mathbb{E}_{P}L(Q^{\prime})$ for every possible $Q^{\prime}$. The log-likelihood $L(P)=-\log p$ is a special case of such a loss function. The risk associated with the log-likelihood identifies the entire distribution $P$ based on the properties of the Kullback–Leibler divergence.
Proof of Theorem 1.
For $h\in\mathcal{H}$ consider the orthogonal decomposition $h=h_{S}+h_{S}^{\perp}$ into the closed subspaces $S_{n}$ and $S_{n}^{\perp}$. From the reproducing property it follows that $h(O_{i})=\langle h_{S}+h_{S}^{\perp}\,,\,K_{O_{i}}\rangle=h_{S}(O_{i})$ and $h_{S}^{\perp}(O_{i})=0$ for every $i\in[n]$. This means that inputs $h$ and $h_{S}$ yield the same value of the empirical risk $\sum_{i}\tilde{L}(O_{i})$ in the objective function of (8). Furthermore, we have assumed that if $h$ is admissible in the optimization problem (8), then so is $h_{S}$. The claim follows by noting that $\lVert h\rVert^{2}_{\mathcal{H}}=\lVert h_{S}\rVert^{2}_{\mathcal{H}}+\lVert h_{S}^{\perp}\rVert^{2}_{\mathcal{H}}$. ∎
C.3 Proof of Proposition 2
Remark C.1.
Assumption 3 assumes that our algorithm converges and terminates in finite time. It is important to highlight that this assumption is commonly adopted in the TMLE literature [29]. As demonstrated in Table D.1, our algorithm, Algorithm 2, converges effectively in practice, typically requiring only a few iterations. While the regularization of the perturbation size (i.e., the norm of $h$) could potentially ensure finite convergence, it is beyond the scope of this paper and remains a topic for future research.
Proof of Proposition 2.
This follows from the first order condition to (8) which must hold at the interior local minimum. First, consider the variations at $p^{\infty}_{n}$ in ~$\widetilde{\mathcal{M}}_{P^{\infty}_{n}}$ along the $n$ directions (scores) given by $K_{P^{\infty}_{n}}({O_{1}},\cdot),\ldots,K_{P^{\infty}_{n}}({O_{n}},\cdot)$ and observe:
$0=$ $\frac{\partial}{\partial_{\alpha_{j}}}\Big{|}_{\alpha=0}\sum_{i}-\log\Big{(}1+\sum_{j}\alpha_{j}(K_{P^{\infty}_{n}})(O_{j},O_{i})\Big{)}p^{\infty}_{n}(O_{i})+\lambda_{K}\alpha^{\top}K_{\mathbf{O}}\alpha$
$\qquad=\left[-\sum_{i}\frac{K_{P^{\infty}_{n}}(O_{j},O_{i})p^{\infty}_{n}(O_{i})}{(1+\sum_{j}\alpha_{j}(K_{P^{\infty}_{n}})(O_{j},O_{i}))p^{\infty}_{n}(O_{i})}+\lambda_{K}(K_{\mathbf{O}}\alpha)_{j}\right]\Bigg{|}_{\alpha=0}$
$\qquad=-\sum_{i}K_{P^{\infty}_{n}}(O_{j},O_{i}).$
This shows that the estimating equation holds for each of these scores, and therefore their linear combinations. Second, by Theorem 1, any $h\in\mathcal{H}_{P^{\infty}_{n}}$ that is orthogonal to these scores must have $h(O_{i})=0$ for all $i\in[n]$ and therefore satisfy Equation (11). From these two observations, we conclude the proposition. ∎
C.4 Proof of Theorem 2
Remark C.2.
Assumptions 2-3 and conclusions derived from them are standard in the TMLE literature [29, 19]. Assumption 4 is mostly technical and ensures that the set of scores that satisfy the estimating equation (11) is sufficiently rich to achieve debiasing for the target parameter by approximating its efficient influence curve estimating equation asymptotically.
Assumption 5 is our main regularity assumption on the model $\mathcal{M}$ and functional $\psi$ needed to control the plug-in bias term $\mathbb{P}_{n}\tilde{\psi}_{\widehat{P}}$ of the proposed KDPE estimator. While this assumption provides a clear intuition for the debiasing mechanism of KDPE, it is not strictly required for the method to work. Less extreme conditions on the regularity of $\mathcal{F}$, such as $C^{\alpha}$ or Sobolev norm bounds (see van der Vaart [52, Ch 19]) and RKHS approximation theory [49]), can imply the Donsker property in S5.
The Donsker assumption S2 has been successfully relaxed for related estimators in the literature [56, 19, 8] via sample-splitting. Our initial analysis of DKPE crucially relies on the Donsker assumption in the argument for asymptotic debiasing (based on S5). We leave the analysis of KDPE with sample-splitting (as done in [56]) to future work.
Proof of Theorem 2.
[…]
Let $h_{j}(\widehat{P})$ be as assumed in S5, then $\text{rem}(j,\widehat{P})\coloneqq P^{*}[\tilde{\psi}_{\widehat{P}}-h_{j}(\widehat{P})]$ can be made arbitrarily small by the assumed uniform approximation and in particular $o_{P^{*}}(1/\sqrt{n})$ for a suitable choice of $j=j(n)$. From S3, and the Donsker and convergence assumptions in S5 we obtain
$\mathbb{P}_{n}\tilde{\psi}_{\widehat{P}}$ $=\mathbb{P}_{n}[\tilde{\psi}_{\widehat{P}}-h_{j}(\widehat{P})]=(\mathbb{P}_{n}-P^{*})[\tilde{\psi}_{\widehat{P}}-h_{j(n)}(\widehat{P})]+\text{rem}(j(n),\widehat{P})=o_{P^{*}}(1/\sqrt{n}).$
We conclude from S2-S5 it follows that $\psi(\widehat{P})$ is asymptotically linear with influence function $\tilde{\psi}_{P^{*}}$:
$\psi(\widehat{P})-\psi(P^{*})=\mathbb{P}_{n}\tilde{\psi}_{P^{*}}-o_{P^{*}}(1/\sqrt{n})+o_{P^{*}}(1/\sqrt{n})+o_{P^{*}}(1/\sqrt{n}).$
Appendix D Simulation Details and Additional Empirical Results
This section provides pseudocode for the bootstrap procedure, empirical justification for our choice of hyperparameters $c,\gamma$, and additional experiments for the bootstrap variance estimation procedure. All simulations were performed on a Dell Desktop with 11th Gen Intel Core i7, 16 GB of RAM.
D.1 Baseline Methods
Initializing $P_{n}^{0}$ for KDPE
For all simulations, we use the SuperLearner [41] and tmle [14] packages in R. For SuperLearner, we set our initial base learners as SL.randomForest, SL.glm, and SL.mean. For tmle, we use the default settings, except for setting the parameter g.bound to 0. For each run of the simulations, we initialize $P_{0}^{n}$ for both TMLE and KDPE to the same initial density estimate to capture the differences in the two debiasing approaches.
Optimization Step Details
For the optimization solver, we use the package CVXR [12] with the default settings to obtain $\alpha^{i+1}$, which corresponds to the parametric update step in Algorithm 2. Because the log-likelihood loss function $L(\alpha,P)=\mathbb{P}_{n}\log((1+\sum_{i=1}^{n}\alpha_{i}K_{0}(o_{i},\cdot))p)$ is concave and all constraints are linear, our update step can be solved by existing convex optimization software. We note that in our experiments, approximately 10% of the runs fail to produce a value using CVXR with default settings due to numerical stability issues. This highlights the potential benefits of alternative loss functions that have known closed-form solutions, such as the mean squared error (MSE), rather than using the log-likelihood loss.
Construction of $K_{P}(o_{i},\cdot)$
To construct our $n$-dimensional parametric submodels, we use the mean-zero RBF kernel $K_{P}(\cdot,\cdot):\mathcal{O}\times\mathcal{O}\rightarrow\mathbb{R}$, which takes the following form:
[…]
where $O,O^{\prime}\in[0,1]\times\{0,1\}^{2}$. Following classical examples [27] and existing software [14] for TMLE, we fix $X\sim\mathbb{P}_{n}(x)$ and $A\sim P_{n}^{0}(A|X)$ (the initial estimate for the propensity score function). This enables us to directly calculate the integral, rather than approximate $K_{P}$:
$K_{P}(O,O^{\prime})=\exp(-\|O-O^{\prime}\|^{2}_{2})-\frac{f_{P}(O)f_{P}(O^{\prime})}{c_{P}},$ (13)
where $f_{P}$ and $c_{P}$ are defined as follows:
$f_{P}(O^{\prime})$ $=f_{P}((x^{\prime},a^{\prime},y^{\prime}))$
$=\frac{1}{n}\sum_{i=1}^{n}\sum_{(a,y)\in\{0,1\}^{2}}P(A=a,Y=y|X=x_{i})\exp(-\|(x^{\prime},a^{\prime},y^{\prime})-(x_{i},a,y)\|^{2}),$
$c_{P}$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{a\in\{0,1\}}\sum_{y\in\{0,1\}}P(A=a,Y=y|X=x_{i})f(x_{i},a,y).$
Note that this calculation is inherent to our setting of binary treatments $A$ and outcomes $Y$ with the empirical marginal distribution for $X$ – other set-ups may not permit a direct calculation of the integral, and may require integral approximations. We defer those experiments to future work.
D.2 Hyperparameter Testing for KDPE
In our experiments, we fixed the sample size at $n=300$. We conducted tests with different values for the density bound $c$, with $c\in\{0.05,0.01,0.002\}$, and for the stopping criteria, we tested $\gamma\in\{0.025,0.005,0.001\}$. For each combination of settings, we performed 100 simulations to obtain the results presented in Table D.1 and Figure 1.
$c$ $\gamma$ Fail % ATE Bias RR Bias OR Bias Avg. Iterations
0.050 0.025 0.05 0.019 0.077 0.205 1.062
0.010 0.025 0.05 0.016 0.063 0.180 1.135
0.002 0.025 0.10 0.026 0.101 0.260 1.120
0.050 0.005 0.09 0.015 0.057 0.150 2.293
0.010 0.005 0.09 0.004 0.026 0.091 2.445
0.002 0.005 0.12 0.006 0.033 0.090 2.505
0.050 0.001 0.23 0.001 0.023 0.068 4.756
0.010 0.001 0.10 0.009 0.040 0.128 5.329
0.002 0.001 0.15 0.009 0.033 0.105 5.802
Table D.1: Average failure percentage, estimate bias, and iterations until convergence.
Both Table D.1 and Figure 1 provide the motivation to use $(c,\gamma)=(0.002,0.001)$. The empirical results shown in Figure 1 coincide with the theoretical results shown in Appendix C, which are derived under the setting where we have no explicit density bound $(c=0)$ and we terminate at a fixed point $(\gamma=0)$. As we see in Figure 1, the simulated distribution of $\psi_{\mathrm{ATE}}(\widehat{P}_{\text{KDPE}})$ with the smallest tested values of $c,\gamma$ obtains a distribution most similar to limiting normal distribution. When the density bound is set to large values (e.g., $c=0.05$), the resulting distributions become skewed and deviate significantly from the normal distribution. This occurs because termination at the boundary (i.e., is tight with respect to the density bound) prevents the distributions from satisfying the gradient conditions shown in Appendix C. When using large values for the convergence tolerance (e.g., $\gamma=0.025$), the algorithm terminates prematurely at a distribution that deviates significantly from the fixed point $P_{n}^{\infty}$. Thus, to maintain the guarantees shown in Appendix C for Algorithm 2, we want the values of $(\gamma,c)$ to be as small as possible. For Sections 4.3 and 4.4, we opt for $c=0.002,\ \gamma=0.001$, the smallest tested hyperparameter values.
The results of this simulation naturally pose questions of how to set $c,\gamma$ as functions of $n$, such that the residual bias caused by these hyperparameters is $o_{P^{*}}(1/\sqrt{n})$. This remains an open question for future work.
D.3 Bootstrap Algorithm
We use the classical bootstrap procedure [21], i.e., sampling $n$ observations from our observed data $\{o_{i}\}_{i=1}^{n}$ with replacement, to estimate the variance of the estimator. We formally state this procedure in Algorithm D.1.
Algorithm D.1 Bootstrap Confidence Interval for KDPE
Input: $\{o_{i}\}_{i=1}^{n}$, convergence tolerance function $l(\cdot)$, convergence tolerance $\gamma$, density bound $c$, and bootstrap iterations $m$.
Obtain distribution $\widehat{P}_{\text{KDPE}}=\text{KDPE}(\{o_{i}\}_{i=1}^{n},l(\cdot),\gamma,c)$.
Set the estimate $\widehat{\psi}=\psi(\widehat{P}_{\text{KDPE}})$.
for $j\in\{1,...,m\}$ do
     Sample with replacement $\{o_{i}^{*}\}_{i=1}^{n}$ from sample $\{o_{i}\}_{i=1}^{n}$.
     Set the estimate $\widehat{\psi}_{j}=\psi(\text{KDPE}(\{o_{i}^{*}\}_{i=1}^{n},l(\cdot),\gamma,c))$
end for
Calculate the mean estimate $\bar{\psi}=\sum_{j=1}^{n}\widehat{\psi}_{j}/n.$
Estimate the variance $\widehat{S}=\sum_{j=1}^{m}(\widehat{\psi}_{j}-\bar{\psi})^{2}/(n-1)$.
Return $\left[\widehat{\psi}\pm\Phi^{-1}(1-\alpha)\sqrt{\widehat{S}}\;\right]$.
D.4 Additional Empirical Results for Bootstrap Confidence Intervals
Building on Section 4.4, we provide additional results regarding the performance of the bootstrap algorithm, Algorithm D.1. In particular, we examine how $m$, the number of bootstrap samples, affects $\widehat{S}$, the estimate of $\text{Var}(\psi(\widehat{P}_{\text{KDPE}}))$, for each of the target parameters. Table D.2 illustrates the bootstrap-estimated variance for KDPE as a function of $m$ and the TMLE variance estimate using a fixed dataset of $(x_{i},a_{i},y_{i})_{i=1}^{n}$.
Method $m$ $\widehat{S}(\widehat{\psi}_{\text{ATE}})$ $\widehat{S}(\widehat{\psi}_{\text{RR}})$ $\widehat{S}(\widehat{\psi}_{\text{OR}})$
KDPE 50 0.00389 0.02841 0.14247
($c=0.01$, 100 0.00473 0.03259 0.15599
$\gamma=0.001$) 150 0.00478 0.03269 0.16241
200 0.00543 0.03704 0.17547
250 0.00519 0.03654 0.17267
KDPE 50 0.00508 0.05384 0.32545
($c=0.002$, 100 0.00530 0.05256 0.30544
$\gamma=0.001$) 150 0.00480 0.05306 0.31699
200 0.00470 0.05477 0.37046
250 0.00441 0.05365 0.36737
TMLE 0.00220 0.01383 0.06105
$\widehat{V}^{\text{sim}}_{\text{KDPE}}$ 0.00322 0.03099 0.17381
($c=0.01,\gamma=0.001$)
$\widehat{V}^{\text{sim}}_{\text{KDPE}}$ 0.00287 0.02555 0.15322
($c=0.002,\gamma=0.001$)
$\widehat{V}^{\text{sim}}_{\text{TMLE}}$ 0.00296 0.02837 0.14981
Table D.2: Estimated variance of $\widehat{\psi}_{\text{ATE}}$, $\widehat{\psi}_{\text{RR}}$, $\widehat{\psi}_{\text{RR}}$ of KDPE and TMLE on one simulation.
To evaluate the dependence of the number of bootstrap samples on variance estimation, we use one draw of $n=300$ samples from data-generating distribution $P^{*}$ and consider different values of $m$, with $m\in\{50,100,150,200,250\}$. For the hyperparameters, we set $\gamma=0.001$, $l(P,P^{\prime})=\|P-P^{\prime}\|^{2}_{L_{2}(\lambda)}$ (where $\lambda$ is the counting measure), and use the density bounds $c\in\{0.01,0.002\}$.
For comparison, we include the variance estimated by TMLE, which is approximated by taking the expectation of the influence function of the target parameter under the empirical measure. Namely, TMLE estimates the variance of the target parameters as follows: $\widehat{S}(\psi_{\mathrm{ATE}}(\widehat{P}_{\text{TMLE}}^{\mathrm{ATE}}))=\mathbb{P}_{n}(\tilde{\psi}_{\widehat{P}_{\text{TMLE}}^{\mathrm{ATE}}}^{\mathrm{ATE}})$, $\widehat{S}(\psi_{\mathrm{RR}}(\widehat{P}_{\text{TMLE}}^{\mathrm{RR}}))=\mathbb{P}_{n}(\tilde{\psi}_{\widehat{P}_{\text{TMLE}}^{\mathrm{RR}}}^{\mathrm{RR}})$, $\widehat{S}(\psi_{\mathrm{OR}}(\widehat{P}_{\text{TMLE}}^{\mathrm{OR}}))=\mathbb{P}_{n}(\tilde{\psi}_{\widehat{P}_{\text{TMLE}}^{\mathrm{OR}}}^{\mathrm{OR}})$, where $\tilde{\psi}_{P}^{\text{target}}$ represents the influence function of the respective target parameter at distribution $P$.
As baselines, we consider the variance of the target parameter estimates, which we estimate using the simulations conducted in Section 4. Namely, we use $n_{\text{sim}}=450$ draws of $n=300$ from the DGP to obtain estimates $\{{\psi}(\widehat{P}_{i})\}_{i=1}^{n_{\text{sim}}}$. Using these estimates, we denote the sample variance as $\widehat{V}^{\text{sim}}=\frac{1}{n_{\text{sim}}-1}\sum_{i=1}^{n_{\text{sim}}}[\psi(\widehat{P}_{i})-\bar{\psi}(\widehat{P}_{i})]^{2}$, where $\bar{\psi}(\widehat{P}_{i})$ is the average estimate of the target parameter across the simulations.
Effect of $m$ on Bootstrap-Estimated Variances
We see that for KDPE with $c=0.01,\gamma=0.001$, the bootstrap-estimated variances $\widehat{S}$ change significantly with $m$ for the ATE and RR target parameters. For $\psi_{\mathrm{ATE}}$ and $\psi_{\mathrm{RR}}$, we observe an approximate 30% change from $m=50$ to $m=250$. For $\psi_{\mathrm{OR}}$, the bootstrap-estimated variance is less sensitive to $m$, only changing approximately 20%. For all target parameters, we see that $\widehat{S}$ changes minimally after $m=200$, which may imply an ideal $m$ to minimize computation costs.
By contrast, for hyperparameters $c=0.002,\gamma=0.001$, we see that the estimated variance for RR, $\widehat{S}(\widehat{\psi}_{\mathrm{RR}})$, changes minimally across $m$, while the estimated variances for ATE and OR are more sensitive to $m$. Generally, we see that KDPE with $c=0.002,\gamma=0.001$ is less sensitive to $m$ relative to $c=0.01,\gamma=0.001$, as changes in $\widehat{S}$ across different values of $m$ for all target parameters tend to be smaller. We stress that these results follow from a single draw of $n=300$ from $P^{*}$, so these trends are by no means conclusive. Rather, they provide a natural starting point for future investigation.
Comparison with Baselines
We compare the variance estimates of KDPE and TMLE to their baselines in the bottom three rows. For all $m$, the bootstrap-estimated variances for $\psi_{\mathrm{ATE}}(\widehat{P}_{\text{KDPE}})$ and $\psi_{\mathrm{RR}}(\widehat{P}_{\text{KDPE}})$ are larger than their estimated baseline variances, while $\psi_{\mathrm{OR}}(\widehat{P}_{\text{KDPE}})$ (for $c=0.01,\gamma=0.001$) is close to its baseline. By contrast, TMLE’s method of approximating the variance underestimates its baseline variance $\widehat{V}_{\text{TMLE}}^{\text{sim}}$ for all three target parameters. This gives important context to the results provided in Section 4.4, which only shows the difference between the variance estimates of KDPE and TMLE without the dependence on the number of bootstrap samples $m$. In particular, these results suggest that the bootstrap procedure in Algorithm D.1 may overestimate the variance of the estimator $\psi(\widehat{P}_{\text{KDPE}})$, leading to conservative confidence intervals for inference. On the other hand, TMLE’s approach to estimating the variance underestimates the variance of $\psi_{\text{target}}(\widehat{P}_{\text{TMLE}}^{\text{target}})$ for all three parameters, which may result in undercoverage issues for inference in a finite-sample setting.
Future Work
These preliminary results are intended as a starting point for future investigation. While Tables 1 and D.2 imply that our bootstrap-estimated confidence intervals are conservative, more extensive testing is required to empirically validate these results. As noted above, Table D.2 only presents the effect of $m$ on a single fixed draw of $n=300$ from the DGP, and cannot answer questions such as average coverage for confidence intervals. This highlights the need for further testing. A large theoretical gap also remains for confirming that our bootstrap-estimated variances are consistent asymptotically, and providing an analysis for Algorithm D.1 under KDPE is another natural step towards our goal of fully computerized inference. Lastly, as shown in Table D.2 and Algorithm D.1, we only use the bootstrap samples to estimate the variance of our estimator and use a normal approximation to construct confidence intervals. Alternative methods for constructing confidence intervals include directly taking the quantiles of the bootstrap estimates and/or subsampling while bootstrapping, which we do not explore in this project.